OpenStudy (mendicant_bias):

I'm going over some theorems regarding the existence of a unique solution for IVP's. Not really sure what it's asking me to do.

OpenStudy (mendicant_bias):

"Find a member of the family that is a solution of the initial value problem." \[y=c_{1}e^{x}+c_{2}e^{-x}, \ \ \ (-\infty,\infty)\]\[y''+y=0, y(0)=0, y'(0)=1\]

OpenStudy (fibonaccichick666):

I believe that is just asking you to solve the IVP

OpenStudy (fibonaccichick666):

find, c1 and c2. At least that is all that it appears to be

OpenStudy (mendicant_bias):

What about all the member family stuff? Yeah, I'm gonna check the answers in the back to get an idea of exactly what they are asking for, but the problem is moreso that I don't understand the theory of exactly what they are asking. My reading on this always leaves me with a pretty hazy understanding, no matter how much I look at it.

OpenStudy (perl):

it looks like a system of equations

OpenStudy (fibonaccichick666):

The family is the general solution without solved c1 and c2

OpenStudy (fibonaccichick666):

memeber is the specific solution to this IVP

OpenStudy (mendicant_bias):

"The given family of functions is the general solution of the differential equation on the indicated interval."

OpenStudy (fibonaccichick666):

yea, and?

OpenStudy (mendicant_bias):

So it's asking for a specific solution to an IVP when it's speaking about a member of a family of solutions?

OpenStudy (mendicant_bias):

(Was just posting that quote for clarity, not correcting you or anybody)

OpenStudy (fibonaccichick666):

yep, specific sln for the specific family member

OpenStudy (fibonaccichick666):

ah ok

OpenStudy (mendicant_bias):

In the second term of every y value, the exponent is supposed to be negative, right?

OpenStudy (fibonaccichick666):

yes, because you have e^-x

OpenStudy (fibonaccichick666):

but you know how to find the first and second deriv of y correct?

OpenStudy (fibonaccichick666):

do you know how to solve the IVP as well?

OpenStudy (mendicant_bias):

Yeah, I know how to find the derivatives of y. A little confused about perl's result, because it looks like you come up with c_{1}=c_{2}, and c_{1}=-c_{2} for the two eqns, a contradiction.

OpenStudy (fibonaccichick666):

yea, just try it and show me your work.

OpenStudy (fibonaccichick666):

I can tell you if it is correct or not once I see your work

OpenStudy (mendicant_bias):

From which point? The very beginning, the derivative, or what?

OpenStudy (mendicant_bias):

(I'll start from the beginning)

OpenStudy (fibonaccichick666):

from the derivatives,

OpenStudy (fibonaccichick666):

@perl, you typed up the question wrong when you did it.

OpenStudy (perl):

oh sorry

OpenStudy (mendicant_bias):

\[y = c_{1}e^{x}+c_{2}e^{-x}; \ \ \ \frac{dy}{dx}=c_{1}e^{x}-c_{2}e^{-x}.\]\[\frac{d^{2}y}{dx^{2}}=c_{1}e^{x}+c^{2}e^{-x}.\]

OpenStudy (fibonaccichick666):

although it did end up not mattering because you did the derivatives correctly

OpenStudy (fibonaccichick666):

ok, so now, plug and chug

OpenStudy (mendicant_bias):

Plugging in to the original Differential Equation: \[y''+y=0;\]\[\bigg[c_{1}e^{x}+c_{2}e^{-x}\bigg]+\bigg[c_{1}e^{x}+c_{2}e^{-x}\bigg]=0\]

OpenStudy (fibonaccichick666):

ok, so now what does that simplify to?

OpenStudy (mendicant_bias):

Not sure where the squared came from, up above was typo.

OpenStudy (fibonaccichick666):

same here, which is why I deleted it as soon as I posted it lol

OpenStudy (mendicant_bias):

Just two times both terms;\[2c_{1}e^{x}+2c_{2}e^{-x}=0\]

OpenStudy (fibonaccichick666):

is there a type-o in the question?

OpenStudy (mendicant_bias):

Let me check.

ganeshie8 (ganeshie8):

y = c1e^x + c2e^(-x) y(0) = 0 : `0 = c1 + c2` y'(0) = 1 : `1 = c1 - c2`

OpenStudy (mendicant_bias):

Yup, sorry, folks, it's minus. y''-y=0, which does indeed equal zero, not sure about the significance of that result, anyways.

OpenStudy (perl):

y = c1*e^x + c2*e^x y' = c1*e^x - c2*e^x y ' ' = c1 * e^x +c2^e^x 1 = y(0) = c1*e^0 + c2*e^0 0 = y'(0) = c1*e^0 - c2*e^0 1 = c1 + c2 0 = c1 - c2 1 = 2*c1 c1 = 1/2 c2 = -1/2

OpenStudy (fibonaccichick666):

ok so now as ganeshie and perl pointed out, you need to solve the system created by the first derivative and the original y

OpenStudy (mendicant_bias):

Alright, so the general procedure for a problem like this with the given information-if I follow-is to take the necessary derivatives (y^(n)), and evaluate those derivatives at the given initial or boundary conditions to then solve for the constants present in the solution

OpenStudy (fibonaccichick666):

yup, solve for your constants then check using the problem and plugging and chugging

OpenStudy (fibonaccichick666):

same thing when you get to laplace IVPs

OpenStudy (perl):

that typo in the question was significant, it threw me off :)

OpenStudy (fibonaccichick666):

same here, I realized it after I said it

OpenStudy (fibonaccichick666):

because we have zeroes

OpenStudy (perl):

you realized there was a typo because it wasnt satisfying y ' ' + y = 0

OpenStudy (perl):

the given solution y(x) was not satisfying y ' ' + y = 0, etc

OpenStudy (perl):

the question was a contradiction

OpenStudy (fibonaccichick666):

I made the fault of only looking at the first and thinking you then had the derivatives wrong as well. But yea, that's why I realized the type-o. Only solution was trivial doesn't make sense

OpenStudy (perl):

right :)

OpenStudy (mendicant_bias):

Alright, I'm getting lost in talking about this problem as to where I move from here; in the situation that we have where the derivatives have been taken, the appropriate equations have been evaluated given the initial or boundary conditions, and the constant coefficients have been solved for, is that it? Or is there anything else to do.

OpenStudy (fibonaccichick666):

yep that's it. You have done it

OpenStudy (fibonaccichick666):

you can use the original to check your answer

OpenStudy (mendicant_bias):

You guys did it, heh. I'll have to work an example on my own, but yeah, thank you, guys.

OpenStudy (perl):

you can check with wolfram (just as a final check) http://www.wolframalpha.com/input/?i=solve+y%27+%27+-+y+%3D+0%2C+y%280%29%3D0%2C+y+%27+%280%29+%3D+1+

OpenStudy (fibonaccichick666):

you actually did it. If you can find the derivatives

OpenStudy (fibonaccichick666):

then once that I trust you can solve the system

OpenStudy (mendicant_bias):

(It's really tempting to just sort of, uh, put my head down here in the library, and fall asleep, lol...)

OpenStudy (perl):

interestingly, the original question I got y = sin x as a solution http://www.wolframalpha.com/input/?i=solve+y%27+%27+%2B+y+%3D+0%2C+y%280%29%3D0%2C+y+%27+%280%29+%3D+1+

OpenStudy (fibonaccichick666):

lol yea? I get that it's late